Mr. Carl spent $7.75 on breakfast. Afterward, he spent 1/4 of the money left in this wallet at the florist. He then had $8.50 left. How much did he start off with?
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

You have two pieces of wood that will make up two sides of a triangular picture frame. One is 6 in. long and the other is 7 in. long. What is the range of the possible lengths for the third side of the frame? Show your work.
-----
Side 3 has to be greater than the difference and less than the sum of the other 2 sides.
-----
1 < s3 < 13

Answer by Theo(13342)

let x equal the amount he started off with.

x - 7.75 is what he has left after he spent 7.75 for breakfast.

3/4 * (x - 7.75) is the amount he has left after he spent 1/4 * (x - 7.75) at the florist.

he had 8.50 left.

therefore 3/4 * (x - 7.75) = 8.50

simplify to get 3/4 * x - 3/4 * 7.75 = 8.50.

simplify further to get 3/4 * x - 5.8125 = 8.50

add 5.8125 to both sides of the equation to get:

3/4 * x = 14.3125.

multiply both sides of the equation by 4/3 to get:

x = 19.08333333333.

the original amount of money is 19.08333333333 dollars.

that's a weird figure but it works out as shown below.

state with 19.08333333333 dollars.

subtract 7.75 to get 11.333333333 dollars.

take 1/4 * 11.333333333 and subtract it from 11.3333333 to get 8.5 dollars.

it works out, even though the original amount doesn't make a lot of sense since who has 19.08333333333 dollars in his pocket?

it you want to round it to the nearest penny, then the original amount is 19 dollars and 8 cents.